OpenStudy (anonymous):

If you were to rearrange this function: 1000=100(K^(.8)*10^(.2)) how would you solve for "K"

OpenStudy (anonymous):

\[1000=100(K ^{0.8}\times 10^{0.2})\]By rewriting 1000 and 100 as powers of 10, the equation becomes:\[10^{3}=10^{2}\times 10^{0.2} \times K ^{0.8}\]We can then bring 10^2 and 10^(0.2) to the left hand side:\[10^{3}/(10^{2}\times 10^{0.2})=K ^{0.8}\]By the rules of exponentiation/powers (sorry, I don't know the exact name):\[10^{3-2-0.2}=10^{0.8}=K ^{0.8}\]At this stage, it is clear that K=10

OpenStudy (anonymous):

Thank you!